OpenStudy (janu16):

If cosine of x equals 1 over 2, what is sin(x) and tan(x)? Explain your steps in complete sentences.

OpenStudy (janu16):

\[\cos(x)=\frac{ 1 }{ 2 }\]

jhonyy9 (jhonyy9):

so you need to know that cos pi/3 = 1/2 what mean that sin pi/3 = sqrt3 /2 and tan pi /3 = sin pi /3 / cos pi /3 = (sqrt3 /2 / 1/2 = (sqrt3 /2) *(2/1) = sqrt3 hope helped and this is understandably sure right easy for you

OpenStudy (janu16):

thanks

jhonyy9 (jhonyy9):

np welcome was my pleasure and anytime good luck bye bye